OpenStudy (anonymous):

Find an equation that models a hyperbolic lens with a=16 inches and foci that are 40 inches apart. Assume that the center of the hyperbola is the origin and the transverse axis is vertical.

OpenStudy (anonymous):

The formula for the equation of a hyperbola of a vertical transverse axis is \[\frac{ Y^2 }{ a^2 }-\frac{ X^2 }{ b^2 }= 1\]

OpenStudy (anonymous):

Dissecting your problem for the values for that formula you can find the equation by using the formula

OpenStudy (anonymous):

Thank You!!!

OpenStudy (awkwardpanda):

do you remember the answer? @rifel7272